what is the value of x : 8x + 1 = 5?
Answer by Theo(13342)

start with 8x + 1 = 5
subtract 1 from both sides of the equation to get 8x = 4
divide both sides of the equation by 8 to get x = 4/8 = 1/2.

the value of x: 8x + 1 = 5 is 1/2.

the value of x: 8x + 1 = 5 means the value of x such that 8x + 1 = 5.

this can only occur if x is equal to 1/2.

8 * 1/2 + 1 = 4 + 1 = 5.